solve LOGa(8x+5) =LOGa(4x+29)
the a is lower than the word LOG. What does this mean?
Someone pleeze help

Answer by ankor@dixie-net.com(12706)

solve LOGa(8x+5) =LOGa(4x+29)
the a is lower than the word LOG. What does this mean?
a is the base of the log, if it is not specified it is assumed to be 10
:
However, it really doesn't matter what the value of "a" is because;
:
If loga(8x+5) equals loga(4x+29),
then,
(8x+5) = (4x+29)
:
easy to solve
8x + 5 = 4x + 29
8x - 4x = 29 - 5
4x = 24
x = 24/4
x = 6
:
You can substitute 6 for x in the original equation and see:
loga(8(6)+5) = loga(4(6)+29)
loga(48+5) - loga(24+29)
loga(53) = loga(53)